OpenStudy (anonymous):

A particle is thrown vertically upwards with a speed of 50m/s the ratio of the distance travelled in 2nd second to the last second of its flight is?

OpenStudy (anonymous):

T = 2u/g = 100/10 = 10s Dn = u + a/2 (2n-1) D2/D10 = (50 + 5*3)/(50+5*19) = 13/27....Is this Correct

OpenStudy (experimentx):

last second = first second

OpenStudy (anonymous):

i got my mistake..Since it is projected upwards a = -g...i took it as +g..)

OpenStudy (anonymous):

it shuld be 50-15 / (50 + 5*19)